Circular No. 7942

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2002 O5 S. Pravdo and K. Lawrence, Jet Propulsion Laboratory, report the

discovery of a comet (with nuclear condensation of diameter ∼ 6′′ and tail extending 12′′ to the southwest) on NEAT images taken at Haleakala on July 30.25 UT. After the object was placed on the NEO Confirmation Page, confirmation of cometary activity was reported by: R. Fredrick, K. Smalley, and R. Trentman, Louisburg, Kansas, with a 0.75-m reflector (tail 40′′ long in p.a. 130^{o}); P. M. Kilmartin, Mt John Observatory, New Zealand, with a 0.6-m reflector (short tail in ∼ p.a. 130^{o}); and R. H. McNaught and G. J. Garradd, Siding Springs Observatory, Australia, with a 1.0-m reflector (tail 13′′ long in p.a. 125^{o}). Selected observations appear below.

2002 UT α_{2000} δ_{2000} m_{1} Observer
July 30.25205 16^{h}06^{m}03.^{s}62 −5^{o}36′39′′.5 17.2 NEAT
31.21119 16 09 40.61 −4 38 20.8 15.3 Fredrick
31.32130 16 10 06.93 −4 31 06.2 17.1 Kilmartin
31.45306 16 10 36.23 −4 23 14.8 16  McNaught

V1548 AQUILAE D. K. Lynch, S. Mazuk, R. J. Rudy, C. C. Venturini, The Aerospace

Corporation; R.C. Puetter, Center for Astrophysics and Space Science, University of California at San Diego; and R. B. Perry, NASA Langely Research Center, report 0.8–2.5 μm spectroscopy using the Aerospace Corporation Near-Infrared Imaging Spectrograph and the Lick Observatory 3-m telescope on July 17.4 UT (∼ 431 days after discovery): “The nova is well into its nebular phase. He I at 1.0830 μm is by far the strongest line in a spectrum that displays full Paschen and Brackett series of H I. He II lines are present as well but no coronal lines are detected. The emission lines are comparatively narrow for a nova with FWHM of 800-900 km/s.”

COMET C/2002 O4 (HÖNIG) Visual m_{1} estimates: July 28.86 UT, 11.3 (M. Reszelski, Szamotuly,

Poland, 0.41-m reflector); 28.94, 10.6 (M. Lehky, Hradec Kralove, Czech Republic, 0.42-m reflector); 29.91, 10.0 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 29.94, 10.2 (Lehky); 29.94, 9.8 (J. J. Gonzalez, Leon, Spain, 0.20-m reflector); 30.71, 10.0 (K. Yoshimoto, Yamaguchi, Japan, 0.25-m refiector).

2002 July 31 \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT Carl W. Hergenrother